# CAE-Net: Generalized Deepfake Image Detection using Convolution and Attention Mechanisms with Spatial and Frequency Domain Features

### v2

@@ Abstract
Effective deepfake detection tools are becoming increasingly essential to the growing usage of deepfakes in unethical practices. There exists a wide range of deepfake generation techniques, which makes it challenging to develop an accurate universal detection mechanism. The 2025 IEEE Signal Processing Cup (DFWild-Cup competition) provided a diverse dataset of deepfake images containing significant class imbalance. The images in the dataset are generated from multiple deepfake image generators, for training machine learning model(s) to emphasize the generalization of deepfake detection. To this end, we proposed a disjoint set-based multistage training method to address the class imbalance and devised an ensemble-based architecture CAE-Net. Our architecture consists of a convolution- and attention-based ensemble network, and employs three different neural network architectures: EfficientNet, Data-Efficient Image Transformer (DeiT), and ConvNeXt with wavelet transform to capture both local and global features of deepfakes. We visualize the specific regions that these models focus on for classification using Grad-CAM, and empirically demonstrate the effectiveness of these models in grouping real and fake images into cohesive clusters using t-SNE plots. Individually, the EfficientNet B0 architecture has achieved 90.79% accuracy, whereas the ConvNeXt and the DeiT architecture have achieved 89.49% and 89.32% accuracy, respectively. With these networks, our weighted ensemble model achieves an excellent accuracy of 94.63% on the validation dataset of the SP Cup 2025 competition. The equal error rate of 4.72% and the Area Under the ROC curve of 97.37% further confirm the stability of our proposed method. Finally, the robustness of our proposed model against adversarial perturbation attacks is tested as well, showing the inherent defensive properties of the ensemble approach.
keywords:
Deepfake detection , ConvNext , Transformer , DeiT , Wavelet , EfficientNet
[…]
[inst1]organization=Department of Electrical and Electronic Engineering,addressline=BUET, city=Dhaka, postcode=1205, country=Bangladesh
[…]
[inst2]organization=School of Data and Sciences,addressline=BRAC University, city=Dhaka, postcode=1212, country=Bangladesh
@@ 1 Introduction
The combination of multiple deep learning architectures has been shown to perform well on deepfake detection. Gowda and Thillaiarasu [3] showed that ensemble techniques of Convolutional Neural Network (CNN) models (ResNeXt and Xception) achieve higher accuracy on deepfake detection. Wodajo and Atnafu [4] combined VGG-16 for feature extraction and Vision Transformer (ViT) for classification, showing promising results on the DFDC dataset. Wolter et al. [5] proposed a wavelet-packet-based analysis for detecting GAN-generated images, highlighting spatial frequency differences in synthetic content and achieving competitive results using combined architectures of wavelets and CNNs. Ricker et al. [6] proposed the usage of frequency domain features extracted via DCT, and achieved high accuracy rates (97.7% for GAN and 73% for diffusion models), highlighting the utility of frequency-domain artifacts. Patel et al. developed an enhanced deep CNN (D-CNN) architecture for deepfake detection, achieving reasonable accuracy and high generalization capability [7]. Patch-Forensics, a technique analyzing smaller patches to detect local image artifacts, achieves 100% average precision on StyleGAN-generated images using an Xception Block 2 classifier [8]. However, Abdullah et al. [9] highlighted two major limitations of the existing networks arguing that the state-of-the-art models often lack control over content and image quality, making them ineffective at detecting high-quality deepfakes if trained on low-quality fake images, and they also lack adversarial evaluation, where an attacker can exploit knowledge of the defense. Works on adversarial perturbation attacks to fool deep neural network (DNN) based detectors are a challenge in deepfake detection [10, 11]. Training on diverse and adversarially perturbed datasets can increase the robustness of a detection architecture [12]. Lin et al. proposed a CNN-based method for deepfake detection that combines multi-scale convolution with a vision transformer [13]. Their approach features a multi-scale module incorporating dilation convolution and depth-wise separable convolution. This design aims to capture more facial details and identify tampering artifacts at various scales [13]. Note that the ViT has gained significant popularity in recent years for deepfake image detection [14, 15].
@@
Research gap and our contribution. The existing detection methods only work well when the test image resembles the features of the images in the training dataset of the detector. That is, the major issues in the existing detection approaches are i) the lack of generalizability of extracted features and as such, ii) the inability to be up to date with new fake image generation techniques [20]. To address the research gap, the DFWild-Cup competition called for novel approaches that can achieve accurate deepfake detection and provided a diverse dataset (consisting of samples from eight different standard datasets) to help train a generalized deepfake detection algorithm. Such a diverse dataset poses unique challenges to formulating accurate image detection methods. In our proposed solution for generalized deepfake image detection, we have employed a weighted ensemble approach, CAE-Net, which utilizes three different neural network architectures with different strengths: i) an EfficientNet architecture, ii) a Data Efficient Image transformer model, and iii) a ConvNeXt model. More specifically, we consolidate the outputs from these architectures and employ a weighted ensemble mechanism to get the final result. We note that the EfficientNet-based architecture, due to its depthwise separable convolution and squeeze and excitation block, allows learning fine spatial artifacts and enhances channel-wise attention [21]. Besides, the self-attention mechanism and the distillation token introduced in the DeiT model allow learning long-range dependencies across the entire image. More specifically, the EfficientNet model excels at capturing local spatial relationships of an image, while the transformer model can capture global features of an image through the self-attention mechanism [22]. Finally, the ConvNeXt model, combined with wavelet transform as a preprocessing technique, is effective in detecting fine-grained artifacts in deepfake images. We have utilized the “Haar” wavelets due to their notable ability to detect abrupt transitions and edges in the data [23]. A weighted ensemble mechanism of these three networks is used as the final solution, which has enabled us to achieve superior performance compared to the models operating alone. Finally, we test our models against adversarial perturbation attacks, which shows the trade-offs between weighted ensemble and majority voting ensemble methods.
@@
The rest of the paper is organized as follows: Section 2 contains data preprocessing techniques, the description of our CAE-Net architecture, consisting of EfficientNet, DeiT, and ConvNeXt. Section 3 contains the accuracy results of multiple standard models on the dataset, along with the implementation details and other evaluation metrics of our proposed approach. This also contains a comparison of our model’s generalization capabilities against multiple existing works and a defense against adversarial perturbations. Finally, Section 4 draws concluding remarks on the contributions presented in the paper.
@@ 2.2.1 Class Balancing for Model Training
One of the key contributions of our paper is to devise a disjoint set-based multistage training approach to address the class imbalance present in the given dataset. Since the number of fake images is approximately five times the number of real images, we divided the fake images ${}\{\mathcal{F}\}$ into five disjoint subsets $\{\mathcal{F}_{1},\ldots,\mathcal{F}_{5}\}$. Each of the three aforementioned models, denoted as $M_{1},M_{2},M_{3}$, are trained in five stages with these disjoint subsets of fake images. At each stage $i$, the model is trained using the entire real image set $\mathcal{R}$ and one subset of fake images $\mathcal{F}_{i}$, such that the training dataset of stage $i$ is given by:
@@
We emphasize that the same set of real images ${}\{\mathcal{R}\}$ are treated as one class during each stage. The training of each model is initialized with pre-trained weights $\theta_{0}$, obtained from an ImageNet [32]-trained version of the corresponding architecture. The model parameters at each stage are updated iteratively using the dataset $\mathcal{D}_{i}$, through the optimization:
@@
This process continues until the model has been trained on all subsets of fake images, ensuring that the model learns from diverse variations of the fake images while preserving knowledge of real images.
@@ 2.2.2 Data Augmentation and Normalization
As mentioned before, we utilized the same set of real images across five training stages to address the class imbalance. However, this evidently increases the chance of overfitting. To mitigate overfitting and enhance the model’s generalization performance, we implemented four data augmentation techniques: random horizontal flips(with 50% probability), random rotations(within $0^{o}-10^{o}$), color jittering (20% variation in brightness, contrast, saturation and hue), and random resized cropping where it produced $224\times 224$ images, while randomly scaling the crop area between 80% and 120% of the original image size. These augmentations enhanced the diversity of the training dataset, reduced overfitting, and improved the model’s ability to generalize across a wider range of deepfake detection scenarios. It should be noted that such augmentations were used only in training the DeiT architectures. For the ConvNeXt, we employed wavelet transform-based feature extraction from the raw images, and data augmentation techniques would result in incorrect feature representation in the spectral domain. Since all three models were pre-trained using the ImageNet dataset, the training images were normalized using the mean and standard deviation of the ImageNet dataset [32]. The input images were resized accordingly to the specific dimensions for the respective architectures.
@@ 2.3 Wavelet Feature Extraction for ConvNeXt
The wavelet transform is shown to be effective for image feature extraction as it decomposes images into multi-resolution frequency components [33]. The ability to decompose an image into different frequency sub-bands while preserving spatial relationships makes it particularly effective at identifying edges, textures, and subtle patterns of an image [34]. Using 2D Wavelet transform, an image can be decomposed into approximate and detail coefficients [23]. As such, we performed a single-level 2D Discrete Wavelet Transform (2D-DWT) using Haar wavelets to produce the approximate and detailed coefficients of the images and used the computed coefficients as features.
@@
[…]
@@
Figure 2: Real and Fake feature images after wavelet transform. The top left, top right, bottom left, and bottom right portions indicate the approximate, horizontal detail, vertical detail, and diagonal detail coefficients of the original image, respectively. Different colormaps are applied to the horizontal, vertical, and diagonal detail coefficients, and their contrasts are increased for better visibility.
@@ 2.4.1 EfficientNet
EfficientNet-B0 is a computationally efficient model whose architecture is optimized via neural architecture search (NAS) to maximize accuracy under a FLOP constraint, resulting in a highly efficient baseline that outperforms models with similar computational cost [21]. The core block of the EfficientNet architecture is the Mobile Inverted Bottleneck Convolution (MBConv) block inspired by MobileNetV2 [35]. The MBConv block consists of a 1x1 Convolution, which expands the input channels to increase feature dimensionality. Then, a 3x3 or 5x5 depthwise convolution is applied to capture spatial features efficiently. Squeeze-and-Excitation (SE) recalibrates channel importance by computing a channel-wise attention vector [36]. The SE block acts as a lightweight attention mechanism to adaptively recalibrate each feature channel’s importance [37]. It is responsible for two main operations:
@@
EfficientNet-B0 uses MBConv1 (expansion factor 1) to MBConv6 (expansion factor 6) variants, with kernel sizes of 3x3 or 5x5 depending on the stage. A brief summary of different layers, the output shapes and the number of parameters in the EfficientNet-B0 architecture is shown in Table 1.
@@ 2.4.2 Data Efficient Image Transformer (DeiT)
Unlike conventional Recurrent Neural Networks (RNNs), transformers process all inputs simultaneously, relying on a self-attention mechanism. This allows transformers to learn long-range dependencies and global contexts of data [22]. The data-efficient image transformer, based on the backbone of the vision transformer, has been shown to work well in cases where training data is limited [39]. Here, an image is divided into a grid of smaller patches as a sequence of input tokens ($N$ number of $16\times 16$ patches), and each patch is associated with a vector through linear embedding. After adding positional encoding for each patch, the vectors are fed into the transformer encoder, where the self-attention mechanism helps to learn the relationship among all the patches [40]. Note that the attention mechanism is based on a trainable associative memory of (key, value) pairs. A query vector $\mathbf{Q}$ is matched against key vectors $\{\mathbf{K}\}$ using dot products, scaled by $\sqrt{d}$, and normalized via a softmax function to produce weights. These weights are used to compute a weighted sum of value vectors $\{\mathbf{V}\}$, producing an output matrix as:
@@
Table 2: Summary of the Data Efficient Image Transformer (DeiT) Model Architecture

### v3

@@ Abstract
The spread of deepfakes poses significant security concerns, demanding reliable detection methods. However, diverse generation techniques and class imbalance in datasets create challenges. We propose CAE-Net, a Convolution- and Attention-based weighted Ensemble network combining spatial and frequency-domain features for effective deepfake detection. The architecture integrates EfficientNet, Data-Efficient Image Transformer (DeiT), and ConvNeXt with wavelet features to learn complementary representations. We evaluated CAE-Net on the diverse IEEE Signal Processing Cup 2025 (DF-Wild Cup) dataset, which has a 5:1 fake-to-real class imbalance. To address this, we introduce a multistage disjoint-subset training strategy, sequentially training the model on non-overlapping subsets of the fake class while retaining knowledge across stages. Our approach achieved $94.46\%$ accuracy and a $97.60\%$ AUC, outperforming conventional class-balancing methods. Visualizations confirm the network focuses on meaningful facial regions, and our ensemble design demonstrates robustness against adversarial attacks, positioning CAE-Net as a dependable and generalized deepfake detection framework.
This manuscript has been published in the Journal of Visual Communication and Image Representation by Academic Press.
Keywords: Deepfake detection, ConvNext, Transformer, DeiT, Wavelet, EfficientNet
@@ 1 Introduction
The combination of multiple deep learning architectures has been shown to perform well on deepfake detection. Gowda and Thillaiarasu [3] showed that ensemble techniques of Convolutional Neural Network (CNN) models (ResNeXt [4] and Xception [5]) achieve higher accuracy on deepfake detection. Wodajo and Atnafu [6] combined VGG-16 [7] for feature extraction and Vision Transformer (ViT) [8] for classification, showing promising results on the DFDC dataset. Wolter et al. [9] proposed a wavelet-packet-based analysis for detecting Generative Adversarial Networks (GAN)-generated images, highlighting spatial frequency differences in synthetic content and achieving competitive results using combined architectures of wavelets and CNNs. Ricker et al. [10] proposed the usage of frequency domain features extracted via Discrete Cosine Transform (DCT) [11], and achieved high accuracy rates (97.7% for GAN and 73% for diffusion models), highlighting the utility of frequency-domain artifacts. Patel et al. developed an enhanced deep CNN (D-CNN) architecture for deepfake detection, achieving reasonable accuracy and high generalization capability [12]. Patch-Forensics, a technique analyzing smaller patches to detect local image artifacts, achieves 100% average precision on StyleGAN-generated images using an Xception Block 2 classifier [13]. However, Abdullah et al. [14] highlighted two major limitations of the existing networks, arguing that the state-of-the-art models often lack control over content and image quality, making them ineffective at detecting high-quality deepfakes if trained on low-quality fake images, and they also lack adversarial evaluation, where an attacker can exploit knowledge of the defense. Works on adversarial perturbation attacks to fool deep neural network (DNN) based detectors are a challenge in deepfake detection [15, 16]. Training on diverse and adversarially perturbed datasets can increase the robustness of a detection architecture [17]. Lin et al. proposed a CNN-based method for deepfake detection that combines multi-scale convolution with a vision transformer (ViT) [18]. Their approach features a multi-scale module incorporating dilation convolution and depth-wise separable convolution. This design aims to capture more facial details and identify tampering artifacts at various scales [18]. Note that the ViT has gained significant popularity in recent years for deepfake image detection [19, 20].
@@
Research gap and our contribution. We note that the existing detection methods only work well when the test image resembles the features of the images in the training dataset of the detector. As such, the major issues in the existing detection approaches are i) the lack of generalizability of extracted features, and ii) the inability to be up to date with new fake image generation techniques [25]. Motivated by this limitation, we propose a weighted ensemble-based deepfake detection framework, CAE-Net, which combines convolutional and attention mechanisms to capture both spatial and frequency-domain cues. The model integrates three different neural network architectures with complementary strengths: i) an EfficientNet architecture [26], ii) a Data-Efficient Image Transformer (DeiT) [27], and iii) a ConvNeXt model [28] enhanced with a wavelet-based preprocessing stage. This hybrid configuration, to the best of our knowledge, is the first to incorporate wavelet-domain feature enhancement within a CNN–Transformer ensemble, effectively leveraging local, global, and frequency-domain representations for generalized deepfake image detection. Since training on highly diverse datasets remains difficult for most existing architectures, we employ the IEEE Signal Processing Cup 2025 (DF-Wild Cup) dataset to demonstrate our proposed CAE-Net’s generalization capabilities. This dataset combines eight publicly available deepfake benchmark datasets under the DeepfakeBench framework and exhibits a pronounced class imbalance (5:1 fake-to-real ratio). To address this challenge, we design a disjoint set-based multistage training strategy, where the fake class is divided into five non-overlapping subsets and trained sequentially along with real samples, using weight warm-starting across stages to retain learned knowledge. The final weighted ensemble employs probability-level late fusion to combine the outputs of the three backbones, leading to statistically significant improvements over individual models and conventional ensemble techniques. As we demonstrate in Section 3, the proposed model offers strong classification performance and maintains robustness under adversarial perturbations, confirming the effectiveness and reliability of CAE-Net in generalized deepfake image detection.
@@
The rest of the paper is organized as follows: Section 2 contains data preprocessing techniques, a description of our CAE-Net architecture, consisting of EfficientNet, DeiT, and ConvNeXt. Section 3 contains the classification results of multiple standard models on the dataset, an ablation study on the different ensemble variations, along with the implementation details and other evaluation metrics of our proposed approach. This also contains a comparison of our model’s generalization capabilities against multiple existing works and a defense against adversarial perturbations. Section  4 provides an overall view of the interpretation of the results, the significance of the findings, and the limitations of our study. Finally, Section 5 draws concluding remarks on the contributions presented in the paper.
@@ 2.2.1 Class Balancing for Model Training
One of the key contributions of our paper is to devise a disjoint set-based multistage training approach to address the class imbalance present in the given dataset. Traditional class-balancing techniques (oversampling, SMOTE, class weighting) present significant limitations for our dataset’s extreme imbalance.
Random oversampling: We empirically validated that replicating 42,690 real images to match 219,470 fakes introduce severe overfitting during model training, as models see identical real samples multiple times across epochs, memorizing specific artifacts rather than learning generalizable features.
Synthetic generation (SMOTE, GANs): Generating $\sim$177,000 synthetic real images risks introducing unrealistic artifacts or distribution shift. Moreover, since the task itself is deepfake detection, authentic real images must represent unmanipulated content. Synthetic augmentation could paradoxically teach models to accept fake-like patterns as real.
Class weighting: While computationally efficient, loss reweighting alone does not prevent the model from being overwhelmed by the sheer volume of fake samples during training, leading to bias toward the majority class despite adjusted gradients.
Our disjoint multistage approach addresses these issues by ensuring a balanced exposure where each stage trains on a 1:1 balanced subset (42,690 real + 43,894 fake images). All real images remain genuine and unmanipulated samples, preserving the true distribution of authentic content. This strategy effectively leverages our full dataset (262,160 images) without the artifacts, overfitting, or distribution shifts inherent to conventional balancing techniques applied to such extreme imbalances.
Since the number of fake images is approximately five times the number of real images, we divided the fake images $\{\mathcal{F}\}$ into five disjoint subsets $\{\mathcal{F}_{1},\ldots,\mathcal{F}_{5}\}$. Each of the three aforementioned models, denoted as $M_{1},M_{2},M_{3}$, is trained in five stages with these disjoint subsets of fake images. At each stage $i$, the model is trained using the entire real image set $\mathcal{R}$ and one subset of fake images $\mathcal{F}_{i}$, such that the training dataset of stage $i$ is given by:
@@
We emphasize that the same set of real images $\{\mathcal{R}\}$ is treated as one class during each stage. The training of each model is initialized with pre-trained weights $\theta_{0}$, obtained from an ImageNet [37]-trained version of the corresponding architecture. The model parameters at each stage are updated iteratively using the dataset $\mathcal{D}_{i}$, through the optimization:
@@
This process continues until the model has been trained on all subsets of fake images, ensuring that the model learns from diverse variations of the fake images while preserving knowledge of real images. While training our models across stages, we maintained the hyperparameter settings stated in Table 8.
@@ 2.2.2 Data Augmentation and Normalization
As mentioned before, we utilized the same set of real images across five training stages to address the class imbalance. However, this evidently increases the chance of overfitting. To mitigate overfitting and enhance the model’s generalization performance, we implemented four data augmentation techniques: random horizontal flips(with 50% probability), random rotations(within $0^{o}-10^{o}$), color jittering (20% variation in brightness, contrast, saturation and hue), and random resized cropping where it produced $224\times 224$ images, while randomly scaling the crop area between 80% and 120% of the original image size. These augmentations enhanced the diversity of the training dataset, reduced overfitting, and improved the model’s ability to generalize across a wider range of deepfake detection scenarios. It should be noted that such augmentations were used only in training the DeiT architecture. For EfficientNet, data augmentation didn’t enhance validation accuracy or generalization, likely due to its convolutional design already effectively captures the local features. In contrast, DeiT showed notable gains in validation accuracy with augmentation, as its attention-based architecture thrives on greater data variety to learn global patterns [38]. Therefore, we applied augmentation only to DeiT in the final model. For the ConvNeXt, we employed wavelet transform-based feature extraction from the raw images, and data augmentation techniques would result in incorrect feature representation in the spectral domain. Since all three models were pre-trained using the ImageNet dataset, the training images were normalized using the mean and standard deviation of the ImageNet dataset [37]. The input images were resized accordingly to the specific dimensions for the respective architectures.
@@ 2.3 Wavelet Feature Extraction for ConvNeXt
The wavelet transform is shown to be effective for image feature extraction as it decomposes images into multi-resolution frequency components [39]. The ability to decompose an image into different frequency sub-bands while preserving spatial relationships makes it particularly effective at identifying edges, textures, and subtle patterns of an image [40]. Using the 2D Wavelet transform, an image can be decomposed into approximate and detail coefficients [41]. As such, we performed a single-level 2D Discrete Wavelet Transform (2D-DWT) using Haar wavelets to produce the approximate and detailed coefficients of the images and used the computed coefficients as features.
@@
    $A$ $=\left(L_{\text{row}}\ast\phi\right)\downarrow_{2,\text{rows}}\qquad\text{(Approximation, LL)}$
    $V$ $=\left(L_{\text{row}}\ast\psi\right)\downarrow_{2,\text{rows}}\qquad\text{(Vertical Details, LH)}$
    $H$ $=\left(H_{\text{row}}\ast\phi\right)\downarrow_{2,\text{rows}}\qquad\text{(Horizontal Details, HL)}$
    $D$ $=\left(H_{\text{row}}\ast\psi\right)\downarrow_{2,\text{rows}}\qquad\text{(Diagonal Details, HH)},$
@@
Figure 2: Real and Fake feature images after wavelet transform. The top left, top right, bottom left, and bottom right portions indicate the approximate horizontal detail, vertical detail, and diagonal detail coefficients of the original image, respectively. Different colormaps are applied to the horizontal, vertical, and diagonal detail coefficients, and their contrasts are increased for better visibility.
@@ 2.4.1 EfficientNet
EfficientNet-B0 is a computationally efficient model whose architecture is optimized via neural architecture search (NAS) to maximize accuracy under a FLOP constraint, resulting in a highly efficient baseline that outperforms models with similar computational cost [26]. The core block of the EfficientNet architecture is the Mobile Inverted Bottleneck Convolution (MBConv) block inspired by MobileNetV2 [42]. The MBConv block consists of a 1x1 Convolution, which expands the input channels to increase feature dimensionality. Then, a 3x3 or 5x5 depthwise convolution is applied to capture spatial features efficiently. Squeeze-and-Excitation (SE) recalibrates channel importance by computing a channel-wise attention vector [43]. The SE block acts as a lightweight attention mechanism to adaptively recalibrate each feature channel’s importance [44, 45]. It is responsible for two main operations:
@@
EfficientNet-B0 uses MBConv1 (expansion factor 1) to MBConv6 (expansion factor 6) variants, with kernel sizes of 3x3 or 5x5 depending on the stage. A summary of different layers, the output shapes, and the number of parameters in the EfficientNet-B0 architecture is shown in Table 1.
@@ 2.4.2 Data Efficient Image Transformer (DeiT)
Unlike conventional Recurrent Neural Networks (RNNs), transformers process all inputs simultaneously, relying on a self-attention mechanism [47]. This allows transformers to learn long-range dependencies and global contexts of data [48]. The data-efficient image transformer, based on the backbone of the vision transformer, has been shown to work well in cases where training data is limited [27]. Here, an image is divided into a grid of smaller patches as a sequence of input tokens ($N$ number of $16\times 16$ patches), and each patch is associated with a vector through linear embedding. After adding positional encoding for each patch, the vectors are fed into the transformer encoder, where the self-attention mechanism helps to learn the relationship among all the patches [49]. Note that the attention mechanism is based on a trainable associative memory of (key, value) pairs. A query vector $\mathbf{Q}$ is matched against key vectors $\{\mathbf{K}\}$ using dot products, scaled by $\sqrt{d}$, and normalized via a softmax function to produce weights. These weights are used to compute a weighted sum of value vectors $\{\mathbf{V}\}$, producing an output matrix as:
@@
Table 2: Summary of the data-efficient image transformer (DeiT) model architecture
@@
Figure 3: Detail of the proposed CAE-Net architecture with weighted ensemble framework.